# Imaginarity of quantum channels: Refinement and Alternative

### v1

@@ 3 Alternative framework for the imaginarity of quantum channels
that is $\phi=p_{1}\tilde{\phi}_{1}+p_{2}\tilde{\phi}_{2}$. From $\bm{(C_{4})}$ We can know that
@@
(STEP 3).
From (1) in (STEP 1) and (2) in (STEP 2), we obtain $C(\phi)=p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2})$. Next we need to prove $C(\phi_{1})=C(\tilde{\phi}_{1})$, that is, we need to prove $C(\chi_{\phi_{1}})=C(\chi_{\tilde{\phi}_{1}})$.
First, we specify the range of index $i$ and $j$, let $i=0,1,\cdots,|B|-1$, $j=0,1,\cdots,|B_{1}|-1$. Let $K_{0}^{1}$ satisfies $\langle i|K_{0}^{1}|j\rangle=\delta_{i,j}$. An intuitive expression of $K_{0}^{1}$ is
@@
where $\hat{O}$ denotes zero matrix with dimension $|B_{2}|\times|B_{1}|$.
Let $K_{n}^{2}$ satisfy $\langle j|K_{n}^{2}|i\rangle=\delta_{i,j+n|B_{1}|}$, $n=0,1,\cdots,\left\lceil|B_{2}|/|B_{1}|\right\rceil$. $\left\lceil a\right\rceil$ is the ceil of $a$. We can give an example to make this easier to understand. See Example 1 after this proof.
@@
$\mathcal{K}_{0}^{1}=\bigoplus_{j=0}^{|A|-1}K_{0}^{1}=\begin{pmatrix}K_{0}^{1}&&\\ &\ddots&\\ &&K_{0}^{1}\end{pmatrix},\mathcal{K}_{n}^{2}=\bigoplus_{j=0}^{|A|-1}K_{n}^{2}=\begin{pmatrix}K_{n}^{2}&&\\ &\ddots&\\ &&K_{n}^{2}\end{pmatrix},$
we can verify $(\mathcal{K}_{0}^{1})^{\dagger}\mathcal{K}_{0}^{1}=I_{|B_{1}|}$ and $\sum_{j=0}^{\left\lceil|B_{2}|/|B_{1}|\right\rceil}(\mathcal{K}_{n}^{2})^{\dagger}\mathcal{K}_{n}^{2}=I_{|B|}$. Thus $\Theta_{1}\in\mathcal{RSC}_{AB_{1}AB}$ with $\Theta_{1}(\cdot)=\mathcal{K}_{0}^{1}\cdot(\mathcal{K}_{0}^{1})^{\dagger}$, $\Theta_{2}\in\mathcal{RSC}_{ABAB_{1}}$ with $\Theta_{2}(\cdot)=\sum_{j=0}^{\left\lceil|B_{2}|/|B_{1}|\right\rceil}\mathcal{K}_{n}^{2}\cdot(\mathcal{K}_{n}^{2})^{\dagger}$. We can also obtain $\Theta_{1}(\chi_{\phi_{1}})=\chi_{\tilde{\phi}_{1}}$ and $\Theta_{2}(\chi_{\tilde{\phi}_{1}})=\chi_{\phi_{1}}$.
@@
$C(\chi_{\phi_{1}})=C(\Theta_{2}(\chi_{\tilde{\phi}_{1}}))\leqslant C(\chi_{\tilde{\phi}_{1}})=C(\Theta_{1}(\chi_{\phi_{1}}))\leqslant C(\chi_{\phi_{1}}).$
@@
Up to now, we prove $\bm{(C_{5})}$ holds: $C(\phi)=p_{1}C(\phi_{1})+p_{2}C(\phi_{2})$.
$\bm{(C_{5})}\Rightarrow\bm{(C_{3})}+\bm{(C_{4})}$:
(STEP 4).
For $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, we introduce an auxiliary system $S$ with dimension $N$ and orthonormal basis $\left\{|n\rangle\right\}_{n=0}^{N-1}$. The system $AB$ formed by Choi state and the auxiliary system $S$ form a combined system $ABS$, of which the basis is $\left\{|j\alpha\rangle\otimes|n\rangle\right\}$ and the whole system is initially in the state $\chi_{\phi}\otimes|0\rangle\langle 0|\triangleq\chi_{\phi}^{ABS}$.
@@
$\Theta^{ABS}(\chi_{\phi}^{ABS})=\sum_{n}(M_{n}\otimes U_{n})(\chi_{\phi}\otimes|0\rangle\langle 0|)(M_{n}\otimes U_{n})^{\dagger}=\sum_{n}M_{n}\chi_{\phi}M_{n}^{\dagger}\otimes|n\rangle\langle n|\triangleq\sum_{n}p_{n}\chi_{\phi_{n}}\otimes|n\rangle\langle n|,$
@@
We can know that $\sum_{n}p_{n}=1$, $\phi_{n}\in\mathcal{C}_{AB_{n}}$ with $|B|=\sum_{n}|B_{n}|$. From $\bm{(C_{5})}$, we get
$\begin{cases}C(\chi_{\phi}^{ABS})=C(\chi_{\phi}\otimes|0\rangle\langle 0|)=C(\chi_{\phi}),\\ C(\Theta^{ABS}(\chi_{\phi}^{ABS}))=C(\sum_{n}p_{n}\chi_{\phi_{n}}\otimes|n\rangle\langle n|)=C(\bigoplus_{n}p_{n}\chi_{\phi_{n}})=\sum_{n}p_{n}C(\chi_{\phi_{n}}).\end{cases}$
From $\bm{(C_{2})}$, we can obtain
$C(\chi_{\phi})=C(\chi_{\phi}^{ABS})\geqslant C(\Theta^{ABS}(\chi_{\phi}^{ABS}))=\sum_{n}p_{n}C(\chi_{\phi_{n}}),$
@@
(STEP 5).
For any $\left\{\phi_{m}\right\}\subset\mathcal{C}_{AB}$ and any probability distribution $\left\{p_{m}\right\}$, we construct the similar auxiliary system $S$ with orthonormal basis $\left\{|m\rangle\right\}_{m}$ and define $\chi_{\phi}^{ABS}=\sum_{m}p_{m}\chi_{\phi_{m}}\otimes|m\rangle\langle m|$. Due to $\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}(I_{|A||B|}\otimes|0\rangle\langle m|)=I_{|A||B||S|}$, $\Theta^{ABS}\in\mathcal{RSC}_{ABSA^{\prime}B^{\prime}S}$ with $\Theta^{ABS}(\cdot)=\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)\cdot(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}$. Then
$\Theta^{ABS}(\chi_{\phi}^{ABS})=\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)\chi_{\phi}^{ABS}(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}=\sum_{m}p_{m}\chi_{\phi_{m}}\otimes|0\rangle\langle 0|.$
We know from $\bm{(C_{2})}$ and $\bm{(C_{5})}$ that the following inequality holds:
$\sum_{m}p_{m}C(\chi_{\phi_{m}})=C(\chi_{\phi}^{ABS})\geqslant C(\Theta^{ABS}(\chi_{\phi}^{ABS}))=C(\sum_{m}p_{m}\chi_{\phi_{m}}),$
that is, condition $\bm{(C_{4})}$ holds.
Therefore, the proof is finished. ∎
Example 1.
Suppose $|B_{1}|=2$ and $|B_{2}|=3$, then $i=0,1,\cdots,4$, $j=0,1$, $n=0,1,2$ and
$K_{0}^{2}=\begin{pmatrix}1&0&0&0&0\\ 0&1&0&0&0\end{pmatrix}\triangleq\begin{pmatrix}I_{|B_{1}|}&O_{|B_{1}|}&\tilde{O}\end{pmatrix},$
$K_{1}^{2}=\begin{pmatrix}0&0&1&0&0\\ 0&0&0&1&0\end{pmatrix}\triangleq\begin{pmatrix}O_{|B_{1}|}&I_{|B_{1}|}&\tilde{O}\end{pmatrix},$
$K_{2}^{2}=\begin{pmatrix}0&0&0&0&1\\ 0&0&0&0&0\end{pmatrix}\triangleq\begin{pmatrix}O_{|B_{1}|}&O_{|B_{1}|}&\tilde{I}\end{pmatrix},$
where $\tilde{O}$ denotes the first $(|B_{2}|\text{ mod }|B_{1}|)$ columns of $O_{|B_{1}|}$, same does $\tilde{I}$.
It is often difficult to verify whether a measure has strong monotonicity and convexity directly, which may be easier after using Theorem 1. When verifying whether a quantifier is a proper measure, we can choose verifying $\bm{(C_{1})}$$\bm{(C_{4})}$ or verifying $\bm{(C_{1})}+\bm{(C_{2})}+\bm{(C_{5})}$.
Under the requirement of $\bm{(C_{5})}$, that is, for any $p_{1}$, $p_{2}$ satisfied $p_{1}+p_{2}=1$, any $\phi_{1}\in\mathcal{C}_{AB_{1}}$ and $\phi_{2}\in\mathcal{C}_{AB_{2}}$, let $\phi=p_{1}\phi_{1}\oplus p_{2}\phi_{2}$, $\phi\in\mathcal{C}_{AB}$ with $|B|=|B_{1}|+|B_{2}|$. We have $\chi_{\phi_{1}}$ and $\chi_{\phi_{2}}$, and we define $\chi_{\tilde{\phi}_{1}}$ and $\chi_{\tilde{\phi}_{2}}$ below.
$\begin{cases}\chi_{\phi_{1}}=\sum_{j,k}|j\rangle\langle k|\otimes\phi_{1}(|j\rangle\langle k|)/|A|,\\ \chi_{\phi_{2}}=\sum_{j,k}|j\rangle\langle k|\otimes\phi_{2}(|j\rangle\langle k|)/|A|,\\ \chi_{\tilde{\phi}_{1}}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|\big{]}/|A|,\\ \chi_{\tilde{\phi}_{2}}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\bm{0}_{1}(|j\rangle\langle k|)\oplus\phi_{2}(|j\rangle\langle k|\big{]}/|A|.\end{cases}$
Thus the following property holds.
@@
The proof of Property 4 is obvious, here is a helpful example.
Example 2.
The following matrix transformations hold and they are invertible.
$\begin{pmatrix}a_{11}&a_{12}&&&&\\ a_{21}&a_{22}&&&&\\ &&a_{33}&&&\\ &&&b_{11}&b_{12}&\\ &&&b_{21}&b_{22}&\\ &&&&&b_{33}\end{pmatrix}\xrightarrow{K_{1}(\cdot)K_{1}^{\dagger}}\begin{pmatrix}a_{11}&&a_{12}&&&\\ &a_{33}&&&&\\ a_{21}&&a_{22}&&&\\ &&&b_{22}&b_{21}&\\ &&&b_{12}&b_{11}&\\ &&&&&b_{33}\end{pmatrix}\xrightarrow{K_{2}(\cdot)K_{2}^{\dagger}}\begin{pmatrix}a_{11}&&a_{12}&&&\\ &b_{11}&&b_{12}&&\\ a_{21}&&a_{22}&&&\\ &b_{21}&&b_{22}&&\\ &&&&a_{33}&\\ &&&&&b_{33}\\ \end{pmatrix},$
where
$K_{1}=\begin{pmatrix}1&&&&&\\ &&1&&&\\ &1&&&&\\ &&&&1&\\ &&&1&&\\ &&&&&1\\ \end{pmatrix},K_{2}=\begin{pmatrix}1&&&&&\\ &&&&1&\\ &&1&&&\\ &&&1&&\\ &1&&&&\\ &&&&&1\\ \end{pmatrix}.$
Due to $K^{\dagger}_{j}K_{j}=I$, then the invertible $\Theta_{j}\in\mathcal{RSC}$ with $\Theta_{j}(\cdot)=K_{j}\cdot K_{j}^{\dagger}$, $j=1,2$.
Just like $\Theta_{1}$ and $\Theta_{2}$ in this example, we can construct multiple invertible channels such that
$\Theta_{1}(\Theta_{2}(\cdots\Theta_{k}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})\cdots))=p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}}.$
In this way, we can know that the equation (3) holds from the condition $\bm{(C_{2})}$.

### v2

@@ 3 Alternative framework for the imaginarity of quantum channels
that is $\phi=p_{1}\tilde{\phi}_{1}+p_{2}\tilde{\phi}_{2}$. From $\bm{(C_{4})}$, it follows that
@@
STEP 3, obtain $\bm{(C_{5})}$ since $\bm{C(\phi_{j})=C(\tilde{\phi}_{j})}$, $\bm{j=1,2}$.
From inequalities (1) and (2) above, we obtain $C(\phi)=p_{1}C(\tilde{\phi}_{1})+p_{2}C(\tilde{\phi}_{2})$. Next, we need to prove that $C(\phi_{1})=C(\tilde{\phi}_{1})$, which requires us to show that $C(\chi_{\phi_{1}})=C(\chi_{\tilde{\phi}_{1}})$.
First, we specify the range of indices $i$ and $j$, let $i=0,1,\cdots,|B|-1$, $j=0,1,\cdots,|B_{1}|-1$. Let $K_{0}^{1}$ satisfies $\langle i|K_{0}^{1}|j\rangle=\delta_{i,j}$. An intuitive expression for $K_{0}^{1}$ is
@@
where $\hat{O}$ denotes the zero matrix with dimension $|B_{2}|\times|B_{1}|$.
Let $K_{n}^{2}$ satisfy $\langle j|K_{n}^{2}|i\rangle=\delta_{i,j+n|B_{1}|}$ for $n=0,1,\cdots,$ $\left\lceil|B_{2}|/|B_{1}|\right\rceil$, where $\left\lceil a\right\rceil$ means the ceil of $a$.
@@
$\mathcal{K}_{0}^{1}=\bigoplus_{j=0}^{|A|-1}K_{0}^{1}=\begin{pmatrix}K_{0}^{1}&&\\ &\ddots&\\ &&K_{0}^{1}\end{pmatrix},$
$\mathcal{K}_{n}^{2}=\bigoplus_{j=0}^{|A|-1}K_{n}^{2}=\begin{pmatrix}K_{n}^{2}&&\\ &\ddots&\\ &&K_{n}^{2}\end{pmatrix}.$
We have $(\mathcal{K}_{0}^{1})^{\dagger}\mathcal{K}_{0}^{1}=I_{|B_{1}|}$ and $\sum_{j=0}^{\left\lceil|B_{2}|/|B_{1}|\right\rceil}(\mathcal{K}_{n}^{2})^{\dagger}\mathcal{K}_{n}^{2}=I_{|B|}$. Thus $\Theta_{1}\in\mathcal{RSC}_{AB_{1}AB}$ with $\Theta_{1}(\cdot)=\mathcal{K}_{0}^{1}\cdot(\mathcal{K}_{0}^{1})^{\dagger}$, $\Theta_{2}\in\mathcal{RSC}_{ABAB_{1}}$ with $\Theta_{2}(\cdot)=\sum_{j=0}^{\left\lceil|B_{2}|/|B_{1}|\right\rceil}\mathcal{K}_{n}^{2}\cdot(\mathcal{K}_{n}^{2})^{\dagger}$. It can be seen that $\Theta_{1}(\chi_{\phi_{1}})=\chi_{\tilde{\phi}_{1}}$ and $\Theta_{2}(\chi_{\tilde{\phi}_{1}})=\chi_{\phi_{1}}$.
@@
$C(\chi_{\phi_{1}})$
$=$ $C(\Theta_{2}(\chi_{\tilde{\phi}_{1}}))$
$\leqslant$ $C(\chi_{\tilde{\phi}_{1}})$
$=$ $C(\Theta_{1}(\chi_{\phi_{1}}))$
$\leqslant$ $C(\chi_{\phi_{1}}).$
@@
Up to now, we prove $\bm{(C_{5})}$ holds: $C(\phi)=p_{1}C(\phi_{1})+p_{2}C(\phi_{2})$. ∎
Under the requirement of $\bm{(C_{5})}$, we have $\chi_{\phi_{1}}$ and $\chi_{\phi_{2}}$ below.
$\begin{cases}\chi_{\phi_{1}}=\sum_{j,k}|j\rangle\langle k|\otimes\phi_{1}(|j\rangle\langle k|)/|A|,\\ \chi_{\phi_{2}}=\sum_{j,k}|j\rangle\langle k|\otimes\phi_{2}(|j\rangle\langle k|)/|A|.\end{cases}$
We also define $\chi_{\tilde{\phi}_{1}}$ and $\chi_{\tilde{\phi}_{2}}$ below.
$\begin{cases}\chi_{\tilde{\phi}_{1}}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\phi_{1}(|j\rangle\langle k|)\oplus\bm{0}_{2}(|j\rangle\langle k|\big{]}/|A|,\\ \chi_{\tilde{\phi}_{2}}=\sum_{j,k}|j\rangle\langle k|\otimes\big{[}\bm{0}_{1}(|j\rangle\langle k|)\oplus\phi_{2}(|j\rangle\langle k|\big{]}/|A|.\end{cases}$
Thus the following property holds.
Next, we prove the reverse result:
Theorem 2.
$\bm{(C_{5})}$ gives rise to $\bm{(C_{3})}$ and $\bm{(C_{4})}$ when condition $\bm{(C_{1})}$ and $\bm{(C_{2})}$ holds.
Proof.
prove $\bm{(C_{3})}$.
For $\Theta\in\mathcal{RSC}_{ABA^{\prime}B^{\prime}}$ with $\Theta(\cdot)=\sum_{m}M_{m}\cdot M_{m}^{\dagger}$, we introduce an auxiliary system $S$ with dimension $N$ and orthonormal basis $\left\{|n\rangle\right\}_{n=0}^{N-1}$. The system $AB$, formed by Choi state and the auxiliary system $S$, constitutes a combined system $ABS$, whose basis is $\left\{|j\alpha\rangle\otimes|n\rangle\right\}$. The entire system is initially in the state $\chi_{\phi}\otimes|0\rangle\langle 0|\triangleq\chi_{\phi}^{ABS}$.
@@
$\Theta^{ABS}(\chi_{\phi}^{ABS})$
$=$ $\sum_{n}(M_{n}\otimes U_{n})(\chi_{\phi}\otimes|0\rangle\langle 0|)(M_{n}\otimes U_{n})^{\dagger}$
$=$ $\sum_{n}M_{n}\chi_{\phi}M_{n}^{\dagger}\otimes|n\rangle\langle n|$
$=$ $\sum_{n}p_{n}\chi_{\phi_{n}}\otimes|n\rangle\langle n|,$
@@
It can be seen that $\sum_{n}p_{n}=1$, $\phi_{n}\in\mathcal{C}_{AB_{n}}$ with $|B|=\sum_{n}|B_{n}|$. From $\bm{(C_{5})}$, we have
$C(\chi_{\phi}^{ABS})=C(\chi_{\phi}\otimes|0\rangle\langle 0|)=C(\chi_{\phi}),$
and
$C(\Theta^{ABS}(\chi_{\phi}^{ABS}))$
$=$ $C(\sum_{n}p_{n}\chi_{\phi_{n}}\otimes|n\rangle\langle n|)$
$=$ $C(\bigoplus_{n}p_{n}\chi_{\phi_{n}})$
$=$ $\sum_{n}p_{n}C(\chi_{\phi_{n}}).$
From $\bm{(C_{2})}$ and above, we can obtain
$C(\chi_{\phi})\geqslant\sum_{n}p_{n}C(\chi_{\phi_{n}}),$
@@
prove $\bm{(C_{4})}$.
For any $\left\{\phi_{m}\right\}\subset\mathcal{C}_{AB}$ and any probability distribution $\left\{p_{m}\right\}$, we construct a auxiliary system $S$ with orthonormal basis $\left\{|m\rangle\right\}_{m}$ and define $\chi_{\phi}^{ABS}=\sum_{m}p_{m}\chi_{\phi_{m}}\otimes|m\rangle\langle m|$. Notably, we have:
$\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}(I_{|A||B|}\otimes|0\rangle\langle m|)=I_{|A||B||S|}$
and $\Theta^{ABS}\in\mathcal{RSC}_{ABSA^{\prime}B^{\prime}S}$ with $\Theta^{ABS}(\cdot)=\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)\cdot(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}$. Then,
$\Theta^{ABS}(\chi_{\phi}^{ABS})$
$=$ $\sum_{m}(I_{|A||B|}\otimes|0\rangle\langle m|)\chi_{\phi}^{ABS}(I_{|A||B|}\otimes|0\rangle\langle m|)^{\dagger}$
$=$ $\sum_{m}p_{m}\chi_{\phi_{m}}\otimes|0\rangle\langle 0|.$
From $\bm{(C_{2})}$ and $\bm{(C_{5})}$, the following inequality holds:
$\sum_{m}p_{m}C(\chi_{\phi_{m}})$
$=$ $C(\chi_{\phi}^{ABS})$
$\geqslant$ $C(\Theta^{ABS}(\chi_{\phi}^{ABS}))$
$=$ $C(\sum_{m}p_{m}\chi_{\phi_{m}}),$
that is, condition $\bm{(C_{4})}$ holds. ∎
Therefore, from the above two theorems, we can give the Alternative framework for the imaginarity of quantum channels:
Theorem 3.
$\bm{(C_{3})}+\bm{(C_{4})}$ is equivalent to $\bm{(C_{5})}$ when condition $\bm{(C_{1})}$ and $\bm{(C_{2})}$ holds.
When verifying whether a quantifier is an imaginarity measure, we can choose to verify either the conditions $\bm{(C_{1})}$$\bm{(C_{4})}$ or the conditions $\bm{(C_{1})}+\bm{(C_{2})}+\bm{(C_{5})}$.
@@
Proof.
We employ the same idea as in Theorem 1, specifically using $\bm{(C_{2})}$ to establish equivalence. Without loss of generality, we assume $|A|=3$, $p_{1}\phi_{1}(|j\rangle\langle k|)\triangleq X_{jk}$ and $p_{2}\phi_{2}(|j\rangle\langle k|)\triangleq Y_{jk}$. Then,
$p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}}=\begin{pmatrix}X_{11}&X_{12}&X_{13}&&&\\ X_{21}&X_{22}&X_{23}&&&\\ X_{31}&X_{32}&a_{33}&&&\\ &&&Y_{11}&Y_{12}&Y_{13}\\ &&&Y_{21}&Y_{22}&Y_{23}\\ &&&Y_{31}&Y_{32}&Y_{33}\end{pmatrix},$
and
$p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}}=\begin{pmatrix}X_{11}&&X_{12}&&X_{13}&\\ &Y_{11}&&Y_{12}&&Y_{13}\\ X_{21}&&X_{22}&&X_{23}&\\ &Y_{21}&&Y_{22}&&Y_{23}\\ X_{31}&&X_{32}&&X_{33}&\\ &Y_{31}&&Y_{32}&&Y_{33}\end{pmatrix}.$
In this way, we need to find some superchannels between $p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}}$ and $\chi_{\phi}$. We can set
$K_{1}=\begin{pmatrix}I_{|B_{1}|}&&&&&\\ &&&I_{|B_{1}|}&&\\ &&&&I_{|B_{1}|}&\\ &I_{|B_{2}|}&&&&\\ &&I_{|B_{2}|}&&&\\ &&&&&I_{|B_{2}|}\end{pmatrix},$
$K_{2}=\begin{pmatrix}I_{|B_{1}|}&&&&&\\ &I_{|B_{2}|}&&&&\\ &&I_{|B_{1}|}&&&\\ &&&I_{|B_{2}|}&&\\ &&&&I_{|B_{1}|}&\\ &&&&&I_{|B_{2}|}\end{pmatrix}.$
It can be known that $K^{\dagger}_{j}K_{j}=I$, $j=1,2$, and the following transformations,
$p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}}\xrightarrow{K_{1}(\cdot)K_{1}^{\dagger}}N\xrightarrow{K_{2}(\cdot)K_{2}^{\dagger}}p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}},$
where $N$ is
$\begin{pmatrix}X_{11}&&&X_{12}&X_{13}&\\ &Y_{11}&Y_{12}&&&Y_{13}\\ &Y_{21}&Y_{22}&&&Y_{23}\\ X_{21}&&&X_{22}&X_{23}&\\ X_{31}&&&X_{32}&X_{33}&\\ &Y_{31}&Y_{32}&&&Y_{33}\end{pmatrix}.$
We define $\Theta_{j}(\cdot)=K_{j}\cdot K_{j}^{\dagger}$, $j=1,2$. It is clear that $\Theta_{j}\in\mathcal{RSC}$. Then the inequality holds:
$C(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})\geqslant C(p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}}).$
Since $K_{1}$ and $K_{2}$ are invertible, the inverse inequality also holds. Therefore, we have:
$C(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})=C(p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}}).$
When the dimension $|A|$ takes other values, we can similarly construct many invertible real superchannels such that
$\Theta_{1}(\cdots\Theta_{k}(p_{1}\chi_{\phi_{1}}\oplus p_{2}\chi_{\phi_{2}})\cdots)=p_{1}\chi_{\tilde{\phi}_{1}}+p_{2}\chi_{\tilde{\phi}_{2}}.$
In this way, we can conclude the equation (3) holds. ∎